# Imagination Policy: Using Generative Point Cloud Models for Learning Manipulation Policies

### v1

@@ 4.3 Real Robot Experiment
Figure 8: Action inference on Mug-Tree with real-sensor data: (a) the observed real-sensor point cloud and the inferred pick, preplace and place action from Imagination Policy, (b) pick generation, (c) preplace generation, and (d) place generation. The top row shows the generated points with orange color and the bottom row demonstrates the configurations of pick, preplace, and place with the calculated rigid transformations. Please note that we used the point cloud from Franka-Emika Panda gripper to train the model and evaluated it with the Robotiq-85 gripper.
We validated Imagination Policy on a physical robot. We trained a multi-task agent from scratch on 3 tasks using a total of just 30 demonstrations. There was no use of the simulated data or pr-rtaining in this experiment - all demonstrations were performed on the real robot.
Settings. The experiment was performed on a UR5 robot with a Robotiq-85 end effector, as shown in Figure 7a. The workspace was a $48\mathrm{cm}\times 48\mathrm{cm}$ region on a table. There were three RealSense 455 cameras mounted pointing toward the workspace. We split the workspace into two parts to place the object and the placement. The segmented point cloud can be directly obtained by cropping the workspace accordingly. To collect the demonstration, we released the UR5 brakes to push the arm physically and record data of the form (initial observation, pick pose, preplace pose, place pose). The combined point cloud $P_{ab}$ was constructed with segmented points and the poses. [^2] Please note that we used the point cloud from Franka-Emika Panda gripper to train the pick model. During testing, we used MoveIt as our path planner to execute the action sequentially.
@@
Results. We collected 10 human demonstrations of each task. Our model was trained for 200k SGD steps with the same settings as the simulated experiments. We evaluated 15 unseen configurations of each task. The results are reported in Table 3. Visualizations of the captured observation and the generated actions are shown in Figure 8. Videos can be found in supplementary materials. Our failures are mainly caused by the distortion of observations and motion planning errors. For example, the handle of the green mug in Mug-Tree task might disappear due to the sensor noise and calibration, which would result in a place failure.
@@ 5 Conclusion
In this work, we propose Imagination Policy, a multi-task model for manipulation pick and place problems. It brings point cloud generation to key-frame manipulation policy learning, which amortizes action prediction during generation by estimating the drift force of each point. We also analyze the key-frame equivariance of the task and implement it in the model by learning rotation-invariant point features. Imagination Policy demonstrates high sample efficiency and better performance on six challenging RLbench tasks against several strong baselines. Finally, we validate our design choices with a multimodal pick-part dataset.
One limitation of the formulation in this paper is that it relies on segmented point clouds. We believe state-of-the-art segmentation models [43, 44] are sufficient to provide high-quality masks. Additionally, our generation process takes 20 seconds with 1000 steps to finish point cloud generation. Fortunately, a large number of works have studied a range of methods for improving the inference speed of diffusion models [45, 46, 47, 48, 49, 50, 24]. We leave applying these existing techniques to future work. Lastly, this paper assumes a fixed one-to-one correspondence between points in the object point clouds and generated point clouds. However, our pipeline of generation and pose estimation proposed here does not strictly require this. Specifically, one can generate point clouds without a correspondence and then train a point cloud registration model to estimate the transformations.
@@ Acknowledgments
This project were supported in part by NSF 1724257, NSF 1724191, NSF 1763878, NSF 1750649, NSF 2107256, NSF 2134178, NSF 2314182 and NASA 80NSSC19K1474. We would like to thank Jung Yeon Park and Nichols Crawford Taylor for their helpful discussions.

### v2

@@ 4.2 Real Robot Experiment
We validated Imagination Policy on a physical robot. We trained a multi-task agent from scratch on 3 tasks using a total of just 30 demonstrations. There was no use of the simulated data or pre-training in this experiment – all demonstrations were performed on the real robot.
Settings. The experiment was performed on a UR5 robot with a Robotiq-85 end effector, as shown in Figure 6a. The workspace was a $48\mathrm{cm}\times 48\mathrm{cm}$ region on a table. There were three RealSense 455 cameras mounted pointing toward the workspace. We split the workspace into two parts to place the object and the placement. The segmented point cloud was directly obtained by cropping the workspace accordingly. To collect the demonstrations, we released the UR5 brakes to push the arm physically and record data of the form (initial observation, pick pose, preplace pose, place pose). The combined point cloud $P_{ab}$ was constructed with segmented points and the poses. During testing, we used MoveIt as our path planner to execute the action sequentially.
@@
Results. We collected 10 human demonstrations of each task. Our model was trained for 200k SGD steps with the same settings as the simulated experiments. We evaluated 15 unseen configurations of each task. The results are reported in Table 2. Visualizations of the captured observation and the generated actions are shown in Appendix 6.5. Videos can be found in supplementary materials. Our failures are mainly caused by the distortion of observations and motion planning errors. For example, the handle of the green mug in Mug-Tree task might disappear due to sensor noise and calibration, which results in a place failure.
@@ 5 Conclusion
In this work, we propose Imagination Policy, a multi-task model for manipulation pick and place problems. It utilizes point cloud generation for key-frame manipulation policy learning by reasoning about the geometric configuration of the goal state. This process amortizes action prediction during generation by estimating the drift force of each point. We also analyze the key-frame equivariance of the task and implement it in the model by learning rotation-invariant point features. Imagination Policy demonstrates high sample efficiency and superior performance on six challenging RLbench tasks against several strong baselines. Finally, we demonstrate that the method can effectively be used to learn manipulation policies on a physical robot. We test our design choices using an ablation study on a multimodal pick-part dataset in Appendix 6.1.
One limitation of the formulation in this paper is that it relies on segmented point clouds. We believe state-of-the-art segmentation models [44, 45] are sufficient to provide high-quality masks. Additionally, our generation process takes 20 seconds with 1000 steps to finish point cloud generation. Fortunately, a large number of works have studied a range of methods for improving the inference speed of diffusion models [46, 47, 48, 49, 50, 51, 22]. We leave applying these existing techniques to future work. Moreover, this paper mainly focuses on rigid-object manipulation. We add one experiment of articulated object manipulation in Appendix 6.3. Our method might also work well for deformable objects. We will explore articulated objects and deformable objects in future work. Lastly, this paper assumes a fixed one-to-one correspondence between points in the object point clouds and generated point clouds. However, our pipeline of generation and pose estimation proposed here does not strictly require this. Specifically, one can generate point clouds without a correspondence and then train a point cloud registration model to estimate the transformations.
@@ Acknowledgments
This project were supported in part by NSF 1750649, NSF 2107256, NSF 2314182, NSF 2134178, NSF 2409351, and NASA 80NSSC19K1474. Dian Wang was also funded by the JPMorgan Chase PhD fellowship. We would like to thank Jung Yeon Park and Nichols Crawford Taylor for their helpful discussions.